1.Write the equation of the line that passes through (2, 2) and (6, 3) in standard form.
2.Write the equation of the line that passes through (2, 4) and (1, -3) in slope-intercept form.
3.

Answer by Alan3354(69443)

1.Write the equation of the line that passes through (2,2) and (6,3) in standard form.
Use determinants:
|x y 1|
|2 2 1| = 0
|6 3 1|
--------
x*(2-3) - y*(2-6) + 1*(6-12) = 0
-x + 4y - 6 = 0
x - 4y = -6
=====================
2.Write the equation of the line that passes through (2,4) and (1,-3) in slope-intercept form.
Find the slope, m = diffy/diffx
m = -7/-1 = 7
---
Use y= mx + b and either point to find b, the y-intercept
4 = 7*2 + b
b = -10
-----
y = 7x - 10
========================
PS Don't put a space after the comma in the points.